# DBudgetKV: Dynamic Budget in KV Cache Compression for Ensuring Optimal Performance

### v1

@@ Abstract
To alleviate memory burden during inference of large language models (LLMs), numerous studies have focused on compressing the KV cache by exploring aspects such as attention sparsity. However, these techniques often require a pre-defined cache budget; as the optimal budget varies with different input lengths and task types, it limits their practical deployment accepting open-domain instructions. To address this limitation, we propose a new KV cache compression objective: to always ensure the full-cache performance regardless of specific inputs, while maximizing KV cache pruning as much as possible. To achieve this goal, we introduce a novel KV cache compression method dubbed DBudgetKV, which features an attention-based metric to signal when the remaining KV cache is unlikely to match the full-cache performance, then halting the pruning process. Empirical evaluation spanning diverse context lengths, task types, and model sizes suggests that our method achieves lossless KV pruning effectively and robustly, exceeding 25% compression ratio on average. Furthermore, our method is easy to integrate within LLM inference, not only optimizing memory space, but also showing reduced inference time compared to existing methods.
DBudgetKV: Dynamic Budget in KV Cache Compression
for Ensuring Optimal Performance
Xuanfan Nithanks: Equal contributions. Partial work was done during Xuanfan’s internship at WeChat AI.   Liyan Xu11footnotemark: 133footnotemark: 3   Chenyang Lyu  Longyue Wang Mo Yu33footnotemark: 3   Lemao Liu33footnotemark: 3   Fandong Meng33footnotemark: 3   Jie Zhou33footnotemark: 3   Piji Lithanks: Corresponding author. 33footnotemark: 3Pattern Recognition Center, WeChat AI xuanfanni@gmail.comliyanlxu@tencent.comlipiji.pz@gmail.com
@@ 1 Introduction
As most large language models (LLMs) follows autoregressive generation Minaee et al. (2024); Chang et al. (2023); OpenAI (2023), they rely on KV Cache to store intermediate states during inference. Specifically, LLMs access the cached key and value vectors of past tokens during the attention calculation, thereby speeding up inference by re-using these KV cache Vaswani et al. (2017); Ainslie et al. (2023); Shazeer (2019). However, as the model size and the input length increase, the memory required to maintain the KV cache also grows proportionally. Since these vectors are usually stored within GPU memory, managing the KV cache efficiently has become crucial to mitigate the overall memory consumption and computational overhead during LLM inference. For instance, Llama3 8B model Dubey et al. (2024) requires 1 GB of KV cache memory for 2K-token inputs, while its 70B counterpart demands a gigantic memory up to 50GB for 20K tokens.
@@
Consequently, recent works have proposed to effectively reduce KV cache through pruning, leveraging the sparsity of the attention mechanism: certain positions are more pivotal during the generation process, while those less important ones could be pruned with minimal performance degradation; thus, the full KV cache is not always necessary. These observations have led to several optimization methods, such as H2O Zhang et al. (2023), ScissorHands Liu et al. (2023), StreamingLLM Xiao et al. (2024), and FastGen Ge et al. (2024), where they aim to identify and retain only the most salient token positions, discarding less critical ones based on certain pruning criteria. Alongside, other orthogonal paradigms have also been proposed without hard-pruning, such as KVMerger Wang et al. (2024) and D2O Wan et al. (2024) that merge KV vectors, and MLA DeepSeek-AI et al. (2024) that operates attention in the latent space.
@@
While previous pruning methods have successfully demonstrated that a decent size of KV cache could be discarded in practice, there exists one common limitation that has not been addressed yet: these pruning techniques typically require a pre-defined KV cache budget, of which its optimal threshold could vary significantly according to specific tasks or inputs. Hence, in real-world scenarios, one would have to carefully tune many budget thresholds for diverse domains, or set a uniform threshold but likely suffering large degradation on certain tasks. We reckon that neither choice is a satisfactory solution, hindering the widespread adoptation of KV pruning techniques.
To further illustrate, Table 1 shows that when the budget is set to 50%, Llama3-8B-Instruct using H2O achieves over 98% of the full-cache performance on NarrativeQA Kociský et al. (2018), whereas under the same conditions on GSM8K Cobbe et al. (2021), a widely-used math benchmark, the performance drastically falls to less than 42% of the full-cache performance.
@@
To address the aforementioned limitation, in this work, we introduce a new objective for KV cache compression: our method aims to strike towards full-cache performance regardless of specific inputs, while maximizing KV cache pruning as much as possible. Towards this objective, we propose an adaptive pruning method, dubbed DBudgetKV, which Dynamically adjusts KV cache retention per input, without the need of pre-defined memory Budget. Particularly, DBudgetKV automatically identifies input-dependent KV cache pruning that leads to minimal performance degradation. A key implication is that the method tends to realize a higher compression ratio for simpler tasks, while allocating more cache resources for complicated tasks.
@@
Our method features a two-step process: 1) after the prefilling stage, rank all input tokens according to an importance metric; 2) discard the KV cache of each token sequentially, until hitting a stopping criterion. In choosing the importance metric and the stopping criterion, DBudgetKV takes into account two important factors from previous works: i) position bias, where tokens at the beginning and the end are generally more significant than those in the middle Xiao et al. (2024); Jiang et al. (2024b); ii) attention score, where tokens receiving higher scores usually contribute more information to subsequent generation Minaee et al. (2024); Chang et al. (2023); Zhang et al. (2023).
We then propose DBudgetKV of empirical efficacy and low overhead. Concretely, tokens are ranked solely by their positions, with no additional computation involved. The ranking strategy signals high significance to the beginning and the end positions, where the initial tokens are the most important, followed by the remaining tokens in a reversing order. Next in the pruning process, we design an attetion-based metric to determine the stopping condition, such that the pruning halts when this metric signals that the remaining KV cache is unlikely to match the full-cache capacity, thereby fulfilling our objective to approximate full performance with dynamic cache pruning.
More specifically, our attention metric compares the difference between the Frobenius form of the full attention matrix and the reduced matrix by setting pruned positions’ score to zero. When the difference exceeds a universal threshold, the process is halted. Especially, this metric empirically correlates well with the generation performance regardless of inputs, serving as an estimator to bridge the pruning state and the subsequent generation. Furthermore, the entire process is implemented efficiently through PyTorch’s built-in operators, where we show that DBudgetKV not only optimizes memory space, but also acheives time reduction compared to the full-cache inference and three KV pruning methods.
@@
Figure 1: The overall workflow of DBudgetKV in Section 3.2. Initially, tokens are ranked based on their positions, followed by the eviction of the least significant tokens (per layer), whose halting condition is determined by the norm value of the reduced attention matrix. The KV cache for the remaining tokens are then preserved.
@@
Experimental results on 13 datasets varying diverse context lengths and tasks, e.g. mathematical and commonsense reasoning, reading comprehension and coding, demonstrate that DBudgetKV achieves our objective effectively and robustly. The resulting inference is on par or even surpasses the full-cache performance with multiple LLMs of different sizes, including Llama3 Dubey et al. (2024), Qwen2.5 Yang et al. (2024), and Mistral Jiang et al. (2024a). Meanwhile, the dynamic pruning of DBudgetKV allocates relatively high budget on math benchmarks but attains high compression ratio by up to 85% on comprehension tasks. The average budget size reaches 63.7% using Llama3-8B. The overall results suggest that DBudgetKV is able to accomplish lossless KV cache pruning with strong generalizability. Besides, it does not rely on specific model architectures and can coexist with other KV cache optimization techniques. We further conduct more experiments for in-depth analysis and ablation studies, depicting the rationality of our various design choices and hyperparameters.
Our contributions can be summarized below:
@@
  • To the best of our knowledge, we are the first to propose the objective that ensures full-cache performance while maximizing the KV cache pruning, which holds greater practical values and deployment potentials.
@@
  • We introduce DBudgetKV that employs a straightforward two-step process, where its dynamic pruning is designed with minimal degradation to the generation performance.
@@ 2 Related Work
Recent advancements in KV cache compression have emerged to address memory constraints for model inference. Scissorhands Liu et al. (2023) observes the repetitive nature of attention patterns and the persistence of token importance. It then preserves the most important tokens based on attention scores to achieve KV cache compression. H2O Zhang et al. (2023) introduces dynamic eviction policies that strategically balance retention of recent tokens and historically significant "heavy hitters". StreamingLLM Xiao et al. (2024) enables infinite-length sequence processing without fine-tuning by stabilizing attention through landmark tokens. SnapKV Li et al. (2024) enhances compression efficiency through attention score-based selection and clustering of critical KV positions. FastGen Ge et al. (2024) proposes an adaptive KV cache management system that customizes retention strategies per attention head.
However, these methods require setting a fixed memory budget, which makes it difficult to consistently maintain full performance across different domains and datasets, limiting their practical application. In contrast, our method imposes no such constraints, targeting full-cache performance while allowing for dynamic budget size.
@@ 3.1 Objective of KV Cache Compression
As mentioned in Sec. 1, most previous works on KV cache pruning require a pre-defined memory budget beforehand, either by fixing the number of cached positions directly Cai et al. (2024); Li et al. (2024), or retaining positions by a fixed ratio of input length Zhang et al. (2023); Wan et al. (2024). While these approaches are effective in certain circumstances, they may encounter practical challenges when deployed in real-world scenarios.
The biggest issue is that the optimal budgets for LLM inference in real-world scenarios are evidently infeasible to enumerate, as they could vary across tasks and domains, especially for open-domain instructions. For example, mathematical tasks are typically concise and require inference based on all given conditions, thus many positions may logically contribute to the generation, necessitating a relatively higher memory budget. Conversely, for article reading comprehension, LLM may only need a small set of critical positions for the answer generation, featuring a smaller budget in many cases.
As a side effect, inference with previous KV cache pruning is likely to experience performance instability across different inputs, which can be observed by our analysis in Sec. 4.2. Overall, the practical value is thereby diminished.
In this work, we resort to a dynamic paradigm that eliminates the need for manually setting a memory budget. We introduce a new objective for KV cache compression, aiming to automatically reduce the KV cache as much as possible, while fulfilling the condition that the method should always strike towards full-cache performance. In this way, one could utilize such pruning method off-the-shelf, without spending time and resources on tuning usage-specific hyperparameters, which we hope would further advance the research development of KV cache compression techniques.
@@ 3.2 DBudgetKV
Towards our objective, we design a novel ranking-based pruning method applied after the prefilling stage, consisting of two steps. First, all tokens are ranked according to an importance metric per Transformers layer. Then, the system keeps removing the KV cache sequentially, until hitting a stopping criterion. Ultimately, our proposed approach does not need to fix an optimal budget for storing KV cache. Rather, the stopping criterion is tied to the norm of attention matrices that works uniformly regardless of the inputs.
The motivation behind our two-step procedure comes from two phenomena brought by previous works. First, there exists position bias such that positions in the beginning and the end are typically important for the subsequent generation, as observed by StreamingLLM Xiao et al. (2024). Second, attention scores are usually an effective indicator of the token importance, as leveraged by many pruning techniques Liu et al. (2023); Zhang et al. (2023); Li et al. (2024). For our objective, we propose to combine the two important properties in our approach, described in details as follows.
@@
It should be noted that our choice on importance ranking is empirically supported rather than theoretically based. We have experimented other attention-based importance ranking, which are provided in Sec. 4.4. Our position-based strategy is shown both empirically effective and superior on computational overhead.
@@
With the initially ranked KV vectors, our method then sequantially decides whether to continue or terminate pruning, which needs an evaluation metric serving as the stopping criterion. Since we evidently cannot know the precise impact of a KV cache position before the generation, we require a metric that correlates well with the resulting performance change when discarding a position, serving as a bridge to ensure a minimal degradation upon the full KV-cache performance.
Inspired by Devoto et al. (2024), we utilize the Frobenius norm of the attention matrix that fits the requirement well empirically. For a matrix $M\in\mathbb{R}^{m\times n}$, the Frobenius norm of $M$ is the matrix-version of the L2 norm, expressed as $||M||_{F}=\sqrt{\sum_{i=1}^{m}\sum_{j=1}^{n}|M_{i,j}|^{2}}$.
We conduct preliminary experiments to validate its reliability. We employ Llama2-7B and randomly mask a certain portion of the KV cache, setting the corresponding values in the attention matrix $A\in\mathbb{R}^{n\times n}$ to zero. As illustrated in Figure 2, there is a certain degree of positive correlation between its Frobenius norm and the model’s performance, which are shown generalized across different domains.
@@
As the input sequence length $n$ increases, the time and space overhead for norm calculation also grows significantly by $O(n^{2})$. Hence, we next aim to reduce the computational scale from the naive norm calculation. We observe that the Frobenius norm of the last row of $A$ also correlates with the model’s performance; equivalently saying, the attention distribution of the latest token is good enough to serve as an estimator. Formally, we utilize the last $k$ rows (latest $k$ tokens) of $A$ and reduce to a single attention vector $A^{\prime}\in\mathbb{R}^{1\times n}$. The score $s_{j}$ for a position $j\in[1,n]$ in $A^{\prime}$ is:
@@
Finally, for each position in the ranking order, our method compares the difference between the Frobenius norm of the original attention matrix $A$, and the Frobenius norm of the reduced attention vector $A^{\prime}$ of which the corresponding positions are set to 00. We set a universal threshold as 1%, such that when the new norm exceeds 1% difference from the original norm, it hits the stopping criterion. The KV vectors of the remaining positions are then all kept, while the preceding positions would have their KV cache discarded. In this way, our method achieves dynamic KV cache pruning while targeting to maintain the full-cache performance, assisted by the norm metric.

### v2

@@ Abstract
To alleviate memory burden during inference of large language models (LLMs), numerous studies have focused on compressing the KV cache by exploring aspects such as attention sparsity. These techniques are often designed with a pre-defined KV budget; however, as the optimal budget varies by different input lengths and task types, the existence of a fixed budget could result in inconsistent performance accepting inputs of diverse domains. To address this limitation, we propose a new KV cache compression objective: to always ensure the full-cache performance regardless of specific inputs, while maximizing KV cache pruning as much as possible. To achieve this goal, we introduce a novel KV cache compression method dubbed DBudgetKV, which features an attention-based metric to signal when the remaining KV cache is unlikely to match the full-cache performance, then halting the pruning process. Empirical evaluation spanning diverse context lengths, task types, and model sizes suggests that our method achieves lossless KV pruning effectively and robustly, exceeding 25% compression ratio on average. Furthermore, our method is easy to integrate within LLM inference, not only optimizing memory space, but also showing reduced inference time compared to existing methods.
@@ 1 Introduction
Most large language models (LLMs) follows autoregressive generation, where they rely on KV Cache to store intermediate states during inference [1, 2, 3]. Specifically, LLMs access the cached key and value vectors of past tokens during the attention calculation, thereby speeding up inference by re-using these KV cache [4, 5, 6]. However, when the model size and the input length increase, the memory required to maintain the KV cache also grows proportionally. Since these vectors are usually stored within GPU memory, managing the KV cache efficiently has become crucial to mitigate the overall memory consumption and computational overhead during LLM inference. For instance, Llama3 8B model [7] requires 1 GB of KV cache memory for 2K-token inputs, while its 70B counterpart demands a gigantic memory up to 50GB for 20K tokens.
@@
Consequently, recent works have proposed to effectively reduce KV cache, leveraging the sparsity of the attention mechanism: certain positions are more pivotal during the generation process, while those less important ones could be pruned with minimal performance degradation; thus, the full KV cache is not always necessary. These observations have led to several optimization methods, such as H2O [8], ScissorHands [9], StreamingLLM [10], and FastGen [11], where they aim to identify and retain only the most salient token positions, discarding less critical ones based on certain pruning criteria. Apart from pruning, other paradigms have been proposed as well to reduce KV consumption, such as KVMerger [12] and D2O [13] that merge or compress KV vectors instead of hard-pruning.
@@
While these studies have successfully demonstrated that a decent portion of KV cache can be discarded in practice, there remains one common issue that has yet to be addressed: these pruning techniques typically require a pre-defined KV cache budget, of which its optimal threshold could vary significantly according to specific tasks or inputs. In real-world scenarios, one would need to carefully tune budget thresholds optimized for different domains; otherwise, setting a uniform threshold in diverse domains may potentially bring large degradation on certain tasks. We reckon the need for a pre-defined threshold as an overlooked limitation that can lead to unstable and inconsistent performance in practice.
Table 1: With pre-defined cache budgets, KV pruning methods on Llama3 exhibit inconsistent performance across domains (percentage relative to full-cache scores).
Methods GSM8K GPQA CoQA
Budget = 90%
H2O 99.1% 96.9% 100.2%
SLM 96.7% 99.2% 100.2%
SnapKV 93.3% 96.1% 99.9%
Budget = 20%
H2O 3.2% 8.5% 96.5%
SLM 4.3% 11.5% 76.1%
SnapKV 1.7% 3.9% 98.6%
To further illustrate, Table 1 shows that when retaining 90% KV cache, Llama3-8B-Instruct could achieve near full-cache performance across three domains. However, the performance is shown inconsistent with 20% budget; especially on the math benchmark GSM8K [14], all three KV pruning methods drastically fall to less than 5% relative to the full-cache performance.
@@
To address the aforementioned limitation, in this work, we propose a new objective for KV cache compression: our method aims to strike towards full-cache performance regardless of specific inputs, while maximizing KV cache pruning as much as possible. Towards this objective, we propose an adaptive pruning method, dubbed DBudgetKV, which Dynamically adjusts KV cache retention per input, without the need of pre-defined memory Budget. In particular, DBudgetKV automatically identifies input-dependent KV cache pruning that leads to minimal performance degradation. A key implication is that the method tends to realize a higher compression ratio for simpler tasks, while allocating more cache resources for complicated tasks.
@@
Our method features a two-step process: (1) after the prefilling stage, rank all input tokens according to an importance metric; (2) discard the KV cache of each token sequentially, until hitting a stopping criterion. In choosing the importance metric and the stopping criterion, DBudgetKV takes into account two important factors from previous works: i) position bias, where tokens at the beginning and the end are generally more significant than those in the middle [10, 15]; ii) attention score, where tokens receiving higher scores usually contribute more information to subsequent generation [1, 2, 8].
To eliminate the pre-defined KV budget while always targeting full-cache performance, the stopping criterion becomes crucial, such that it should signal whenever the remaining KV cache is unlikely to match the full-cache capacity, thereby fulfilling the objective to approximate full performance with dynamic cache pruning. In DBudgetKV, we design a simple attention-based metric as the stopping condition: the metric compares the difference between the Frobenius form of the full attention matrix and the reduced attention matrix of which pruned positions’ scores are set to zero. When the difference exceeds a universal threshold, the pruning process halts. In practice, we find that this metric empirically correlates well with the full-cache performance regardless of input types, serving as an estimator to bridge the pruning state and the subsequent generation.
@@
Figure 1: The overall workflow of DBudgetKV in Section 3.2. After prefilling, tokens are initially ranked based on their positions, followed by the eviction of the least significant tokens (per layer), whose halting condition is determined by the norm value of the reduced attention matrix. The KV cache for the remaining tokens are then preserved to subsequent generation.
@@
To examine the performance consistency across diverse domains, our experiments adopt 13 datasets varying diverse context lengths and tasks, e.g. mathematical and commonsense reasoning, reading comprehension and coding, demonstrating that DBudgetKV achieves our objective effectively and robustly. With dynamic KV budgets, the resulting inference is on par or even surpasses the full-cache performance with multiple LLMs of different sizes, including Llama3 [7], Qwen2.5 [16], and Mistral [17]. The dynamic pruning of DBudgetKV automatically allocates relatively high cache budget on math benchmarks to prevent degradation, while achieving high compression ratio on comprehension tasks by up to 85%. With Llama3-8B, DBudgetKV reaches an average budget ratio of 63.7% with minimal degradation across inputs. Overall, empirical results suggest that our proposed method is able to accomplish lossless KV cache pruning with strong generalizability. Besides, it does not rely on specific model architectures, and is orthogonal to other KV optimization techniques. We further conduct experiments for in-depth analysis and ablation studies, where we validate that each step of DBudgetKV is necessary in achieving the desired performance, depicting the rationality of various design choices and hyper-parameters. Our contributions can be summarized as follows:
@@
  • To the best of our knowledge, we are the first to call for the objective that ensures full-cache performance while maximizing the KV cache pruning, addressing the limitation of inconsistent performance with fixed KV budget thresholds.
@@
  • We introduce DBudgetKV that employs a straightforward two-step process, where the dynamic pruning is halted with minimal degradation to the generation performance.
@@ 2 Related Work
Recent advancements in KV cache compression have emerged to address memory constraints for model inference. Scissorhands [9] observes the repetitive nature of attention patterns and the persistence of token importance. It then preserves the most important tokens based on attention scores to achieve KV cache compression. H2O [8] introduces dynamic eviction policies that strategically balance retention of recent tokens and historically significant "heavy hitters". StreamingLLM [10] enables infinite-length sequence processing without fine-tuning by stabilizing attention through landmark tokens. SnapKV [18] enhances compression efficiency through attention score-based selection and clustering of critical KV positions. FastGen [11] proposes an adaptive KV cache management system that customizes retention strategies per attention head. Other paradigms beyond KV pruning have been proposed along the way, such as KV cache merging [19].
As prior KV pruning methods typically require setting a pre-defined memory budget, they are by design inherently unable to ensure consistent performance across various domains and datasets, which could prevent their widespread deployment in practice. Our method targets to lift such constraint, achieving near full-cache performance and dynamic budget size without the need of additional tuning.
Attention Pattern
The study of attention head specialization has paved the way for optimizing memory and computation in Transformer models. Early research on encoder architectures shows that some attention heads have persistent positional biases, like consistently focusing on adjacent tokens, indicating redundancy in their learned patterns. Rho-1 [20] reveals that the importance of individual heads varies significantly across tasks, suggesting that not all attention computations are equally crucial during inference, and reveal the potential for dynamically pruning or approximating redundant computations. Recent works [6, 21] have shifted the focus to decoder-only models, where the quadratic memory overhead of retaining intermediate KV vectors poses a significant challenge. Focused Transformer [22] extends the context length of LLMs through a contrastive learning-inspired training process. Landmark-Attention [23] uses a “landmark” token to represent each block of the input and trains the attention to use it for selecting relevant blocks, enabling retrieval of blocks directly through the attention mechanism instead of by relying on a separate mechanism.
Our work leverages attention’s structural regularities, particularly the predictable, position-specific behaviors of certain heads, to dynamically reduce the memory footprint of cached KV states. By identifying and exploiting the properties with position bias and attention distribution, we enable efficient inference without compromising model performance, aligning with the broader goal of dynamic KV cache pruning for various inputs.
@@ 3.1 Objective of KV Cache Compression
As mentioned in Sec. 1, most previous works on KV cache pruning require a pre-defined memory budget beforehand, either by fixing the number of cached positions directly [24, 18], or retaining positions by a fixed ratio of input length [8, 13]. While these approaches are effective in certain circumstances, they may encounter practical challenges when deployed in real-world scenarios.
It is forthcoming that the optimal budgets for LLM inference in real-world scenarios are evidently infeasible to enumerate, as they could vary across tasks and domains, especially for open-domain instructions. For example, mathematical tasks are typically concise and require inference based on all given conditions, thus many positions may logically contribute to the generation, necessitating a relatively higher memory budget. Conversely, for article reading comprehension, LLM may only need a small set of critical positions for the answer generation, featuring a smaller budget.
As a side effect, inference with fixed KV cache pruning may experience performance instability across different inputs, which can be observed by our analysis in Sec. 4.2. We regard it as a practical limitation that motivates the inception of this work.
In addressing this limitation, we resort to a dynamic paradigm that eliminates the need for manually setting a memory budget. We introduce a new objective for KV cache compression, aiming to automatically reduce the KV cache as much as possible, while fulfilling the condition that the method should always strike towards full-cache performance. In this way, one could utilize such pruning method off-the-shelf, without spending time and resources on tuning usage-specific hyper-parameters, which we hope would further advance the research development of KV cache compression techniques.
@@ 3.2 DBudgetKV
Towards our objective, we design a novel ranking-based pruning method applied after the prefilling stage, comprising two steps. First, all tokens are ranked according to an importance metric per Transformers layer. Then, the system keeps removing the KV cache sequentially, until hitting a stopping criterion. Ultimately, our proposed approach does not need to fix an optimal budget for storing KV cache; rather, it relies on the stopping criterion to determine dynamically, which should work uniformly regardless of the inputs.
The motivation behind our two-step procedure comes from two phenomena brought by previous works. First, there exists position bias such that positions in the beginning and the end are typically important for the subsequent generation, as observed by StreamingLLM [10]. Second, attention scores are usually an effective indicator of the token importance, as leveraged by many pruning techniques [9, 8, 18]. For our objective, we propose to combine the two important properties in our approach, described in details as follows.
@@
Like previous KV compression works, our choice on importance ranking is empirically supported rather than theoretically based. We have also experimented other attention-based importance ranking, which are provided in Sec. 4.5. The choice of position-based ranking is empirically shown both empirically effective and superior on computational overhead.
@@
With the initially ranked KV vectors, our method then sequentially decides whether to continue or terminate pruning, which requires an evaluation metric as the stopping criterion. Since one evidently cannot know the precise impact of a KV cache position before the generation, we adopt a metric that empirically correlates well with the performance change when discarding a position, serving as a bridge to ensure a minimal degradation upon the full cache performance.
Inspired by Devoto et al. [25], we utilize the Frobenius norm of the attention matrix that fits the requirement well empirically. For a matrix $M\in\mathbb{R}^{m\times n}$, the Frobenius norm of $M$ is the matrix-version of the L2 norm, expressed as $||M||_{F}=\sqrt{\sum_{i=1}^{m}\sum_{j=1}^{n}|M_{i,j}|^{2}}$.
@@
We conduct preliminary experiments to validate its reliability. We employ Llama2-7B and randomly mask a certain portion of the KV cache, setting the corresponding values in the attention matrix $A\in\mathbb{R}^{n\times n}$ to zero. As illustrated in Figure 2, there is a certain degree of positive correlation between its Frobenius norm and the model’s performance, which are shown generalized across different domains. The simple Frobenius norm works empirically well in subsequent large-scale evaluation in Sec. 4.
As the input sequence length $n$ increases, the time and space overhead for norm calculation also grows significantly by $O(n^{2})$. Hence, we next aim to reduce the computational scale from the naive norm calculation. We observe that the Frobenius norm of the last row of $A$ also correlates with the model’s performance; equivalently speaking, the attention distribution of the latest token is good enough to serve as an estimator. Formally, we utilize the last $k$ rows (latest $k$ tokens) of $A$ and reduce to a single attention vector $A^{\prime}\in\mathbb{R}^{1\times n}$. The score $s_{j}$ for a position $j\in[1,n]$ in $A^{\prime}$ is:
@@
Finally, for each position in the ranking order, DBudgetKV compares the difference between the Frobenius norm of the original attention matrix $A$, and the Frobenius norm of the reduced attention vector $A^{\prime}$ of which the pruned positions are set to 00. We set a universal threshold as 1%, such that when the new norm exceeds 1% difference from the original norm, it hits the stopping criterion. The KV vectors of the remaining positions are then all kept, while the preceding positions have their KV cache discarded. In this way, our method achieves dynamic KV cache pruning while targeting to maintain the full-cache performance, assisted by the norm metric.